OpenStudy (anonymous):

-114=-6(1-6n), what is n?

OpenStudy (anonymous):

distribute the 6 so -144 = -6(36n) then divide by 6 -144/6 = 36n solve -24 = 36n divide by 36 -24/36 = n

OpenStudy (anonymous):

so whats n?

OpenStudy (anonymous):

-2/3 just simplify the fraction

OpenStudy (anonymous):

oh thanks